# When Sharpening Becomes Collapse: Sampling Bias and Semantic Coupling in RL with Verifiable Rewards

### v1

@@ 4.2 Evaluation Result
Comparison with state-of-the-art methods. Table 2 and Table 3 summarize the AVG@8 and PASS@8 performance of different methods, respectively, across six mathematical reasoning benchmarks and three Qwen3 backbones. Overall, our method attains the best AVG@8 and PASS@8 in almost all settings across models and datasets, indicating that calibrated sharpening consistently improves both average quality and success probability. In contrast, different baselines occasionally excel on specific benchmarks. For example, Reinforce++ is competitive on AIME24 and MATH500, and RLOO sometimes attains strong averages. To illustrate this in more detail, consider the Qwen3‑14B results. Here, all baselines are already strong, yet our method still delivers clear gains. In terms of AVG@8, our method can raise the overall average from roughly 70.2 (Reinforce++ / DrGRPO) to 71.15. For PASS@8, the improvement is more striking: from 77.03 (best baseline) to 78.81. In particular, improvements on harder, less saturated benchmarks (AIME24, Minerva, O‑Bench) are more pronounced than on easier ones such as GSM8K, suggesting that our approach is especially beneficial when reasoning requires exploring multiple reasoning modes rather than amplifying a single dominant pattern. On the smaller backbones Qwen3‑4B and Qwen3‑8B, we observe a consistent pattern.
@@
5 Related Work
Since the release of DeepSeek-R1 (DeepSeek-AI, 2025), RLVR (or GRPO and its variants) has gained particular traction for enhancing the reasoning ability of LLMs. To stabilize the often volatile training dynamics, a significant body of work focused on refining advantage estimation. Prominent techniques include RLOO (Leave-One-Out) (Kool et al., 2019), DrGRPO (Liu et al., 2025), which modifies the variance term to improve stability, and Reinforce++ with baseline (Hu et al., ) that shifts from sample-level to global batch-level normalization. Another line of work has identified a critical correlation between model performance and the evolution of policy entropy (Cui et al., 2025). While rapid entropy reduction can accelerate convergence, it drives the model toward premature saturation at local optima. Xiong et al. (2025) highlighted the importance of negative samples in RLVR. Other approaches focus on sustaining entropy to prolong exploration. These methods employ techniques such as safe or flexible clipping mechanisms (Yu et al., 2025; Wu et al., 2025), restricting gradient updates to high-entropy tokens (Wang et al., 2025), or incorporating entropy explicitly as a regularization term (He et al., 2025). However, the field remains divided. Recent empirical studies (Karan & Du, 2025; Yue et al., 2025; Song et al., 2025) indicate that RLVR training can sometimes lead to a degradation in general capabilities.
@@ 6 Conclusion
In this work, we formalised the phenomenon of over-sharpening within RLVR framework. We revealed that the very clarity of verifiable rewards, when combined with finite-sample bias and semantic coupling, can lead to a pathological collapse of reasoning diversity. To counter this, we introduced two calibration mechanisms designed to stabilize the RLVR partition function and decouple the sampling process from the current policy’s modes. We conducted extensive evaluations and showed that these mechanisms substantially mitigate entropy collapse and stabilize generalization in verifiable‑reward settings. Future work may extend our analysis to richer reward structures and more complex interaction protocols (e.g., multi‑turn tool use), etc.
@@
Appendix B Distributional Update Geometry
Theorem B.1.
@@ Appendix B Distributional Update Geometry
The batch-optimal policy $\hat{\pi}$ satisfies the stationary condition $\nabla_{\pi}\hat{J}=0$, which yields $\hat{\pi}(o_{i})=\frac{1}{Z^{\prime}}\pi_{\text{ref}}(o_{i})\exp(\frac{N_{i}A_{i}}{\beta G})$ according to LABEL:lem:batch_opt. Taking the log:
@@
Appendix C Proof of Theorem 2.3
@@
Appendix D Background for Section 2.3
@@
Appendix E Proof of Theorem 2.4
Theorem E.1.
@@
According to 2.2, we have $\mathbf{k}_{q^{\prime}}]_{s}=k((q^{\prime},o^{\prime}),(q,o^{(s)}))\approx\eta K((q,o^{\prime}),(q,o^{(s)}))$.
@@
Appendix F Proof of Corollary 2.2
Corollary F.1.
@@
Appendix G Supplementary Experiment for Section 2.4
@@
Kernel similarity. In Section 2.3, we establish that the propagation of collapse is governed by the ratio $\mathcal{R}$, which is heavily influenced by the ratio $\lambda/\rho$. To quantify this in our toy model, we calculate the dot products of the embeddings to approximate the kernel entries $K_{qq}\approx e_{q}^{\top}e_{q^{\prime}}$. We study three degrees of similarity between the training query ($q=\text{Persian}$) and the held-out query ($q^{\prime}=\text{Siamese}$), corresponding to the High, Mid, and Low similarity settings described in Section 2.4. The fixed embedding for Persian is $e_{Persian}=[0.75,0.5,0.25,0.1]$. The similarity metrics are calculated as follows:
  1. 1.
    High Similarity:
    • $e_{Siamese}=[0.25,0.5,0.75,0.1]$.
[…]
    • Cross-similarity ($\rho\approx e_{Persian}^{\top}e_{Siamese}$): 0.75(0.25)+0.5(0.5)+0.25(0.75)+0.1(0.1)=0.6350.75(0.25)+0.5(0.5)+0.25(0.75)+0.1(0.1)=0.635.
    • Ratio ($\lambda/\rho$): $1.38$.
  2. 2.
    Mid Similarity:
    • $e_{Siamese}=[0.1,0.5,0.9,0.1]$.
    • Cross-similarity: 0.75(0.1)+0.5(0.5)+0.25(0.9)+0.1(0.1)=0.560.75(0.1)+0.5(0.5)+0.25(0.9)+0.1(0.1)=0.56.
    • Ratio ($\lambda/\rho$): $1.57$.
  3. 3.
    Low Similarity:
    • $e_{Siamese}=[0.0,0.5,1.0,0.1]$.
    • Cross-similarity: 0.75(0.0)+0.5(0.5)+0.25(1.0)+0.1(0.1)=0.510.75(0.0)+0.5(0.5)+0.25(1.0)+0.1(0.1)=0.51.
    • Ratio ($\lambda/\rho$): $1.72$.
These calculations demonstrate that as the embeddings become less similar (moving from High to Low), the cross-similarity $\rho$ decreases, causing the ratio $\lambda/\rho$ to increase. A higher $\lambda/\rho$ implies weaker semantic coupling, meaning updates to Persian have less impact on Siamese. This aligns with the empirical results in Figure 2, where the ”High Similarity” setting (lowest $\lambda/\rho$) exhibited the most rapid degradation of accuracy on the held-out class.
Appendix H Supplementary Material for Section 3
@@
Appendix I Supplementary Experiment for Section 4
Table 5: Sensitivity of IAC to the calibration strength $\alpha$. We report AVG@8 and PASS@8 on Math500 and O-Bench for Qwen3-8B when varying $\alpha$ in IAC. When $\alpha=0$, IAC reduces to the vanilla GRPO objective. We also include a simple filtering-based baseline that discards the easiest 25% of training queries according to their empirical success rate.
Metric Dataset 0 0.5 1 1.5 2 Filtering-based
@@ Appendix I Supplementary Experiment for Section 4
Table 5 also compares IAC against a filtering strategy that discards the easiest 25% of training queries according to their empirical success rate. Across both Math500 and O-Bench, filtering yields only modest gains over GRPO and is consistently weaker than IAC. This gap highlights a key limitation of hard filtering: removing easy queries entirely reduces the effective training distribution and discards useful gradient information, particularly about how to maintain correctness on already-solved patterns. IAC, in contrast, keeps all queries in the optimization loop but rescales their influence on the update. Easy queries contribute a smaller—but non-zero—signal, preventing them from overwhelming the batch partition function $Z^{\prime}(q)$, while hard queries are emphasized in a smooth, data-dependent manner. From a distributional perspective, filtering crudely prunes regions of the solution space, whereas IAC gently rebalances probability mass across difficulty levels. The latter is more effective at preserving useful behaviors on easy instances while still allocating sufficient capacity to explore and stabilize difficult, low-success regimes, which explains the consistent empirical advantage of IAC over filtering-based baselines.

### v2

@@ 5.2 Evaluation Result
Comparison with state-of-the-art methods. Table 2 summarizes the AVG@8 and PASS@8 performance of different methods, respectively, across six mathematical reasoning benchmarks and three Qwen3 backbones. Overall, our method attains the best AVG@8 and PASS@8 in almost all settings across models and datasets, indicating that calibrated sharpening consistently improves both average quality and success probability. In contrast, different baselines occasionally excel on specific benchmarks. For example, Reinforce++ is competitive on AIME24 and MATH500, and RLOO sometimes attains strong averages. To illustrate this in more detail, consider the Qwen3‑14B results. Here, all baselines are already strong, yet our method still delivers clear gains. In terms of AVG@8, our method can raise the overall average from roughly 70.2 (Reinforce++ / DrGRPO) to 71.15. For PASS@8, the improvement is more striking: from 77.03 (best baseline) to 78.81. In particular, improvements on harder, less saturated benchmarks (AIME24, Minerva, O‑Bench) are more pronounced than on easier ones such as GSM8K, suggesting that our approach is especially beneficial when reasoning requires exploring multiple reasoning modes rather than amplifying a single dominant pattern. On the smaller backbones Qwen3‑4B and Qwen3‑8B, we observe a consistent pattern.
@@ 6 Conclusion
In this work, we formalized the phenomenon of over-sharpening within the RLVR framework. Our theoretical analysis identifies two fundamental drivers of distribution collapse: finite-sample bias, which causes the uniform suppression of unsampled correct modes, and semantic coupling, which allows this collapse to propagate across related tasks through shared model parameters. To mitigate these issues, we introduced two plug-and-play calibration mechanisms IAC and DLC that stabilize the partition function and diversify the sampling process. Extensive evaluations across multiple reasoning benchmarks and model scales demonstrate that our methods effectively achieve consistent performance gains over state-of-the-art baselines.
@@
Appendix B Proof of Theorem 3.2
Proof.
We begin by re-expressing the empirical objective $\hat{J}(\pi)$ at the mode level. Given a query $q$, the objective is a combination of the sample-weighted advantage and the KL divergence relative to the reference policy:
$\hat{J}(\pi)=\sum_{i}\frac{N_{i}}{G}A(q,o_{i})\log\pi(o_{i}|q)-\beta\sum_{i}\pi(o_{i}|q)\log\frac{\pi(o_{i}|q)}{\pi_{ref}(o_{i}|q)},$
where we treat $\pi$ as a functional optimization target over the discrete output space.
To find the maximizer $\hat{\pi}$, we introduce a Lagrange multiplier $\lambda$ for the normalization constraint $\sum_{i}\pi(o_{i})=1$. The Lagrangian $\mathcal{L}$ is defined as:
$\mathcal{L}(\pi,\lambda)=\sum_{i}\left[\frac{N_{i}A_{i}}{G}\log\pi(o_{i})-\beta\pi(o_{i})\log\frac{\pi(o_{i})}{\pi_{ref}(o_{i})}\right]-\lambda\left(\sum_{i}\pi(o_{i})-1\right).$
Taking the functional derivative with respect to each $\pi(o_{i})$ and setting it to zero yields:
$\frac{\partial\mathcal{L}}{\partial\pi(o_{i})}=\frac{N_{i}A_{i}}{G\cdot\pi(o_{i})}-\beta\left(\log\frac{\pi(o_{i})}{\pi_{ref}(o_{i})}+1\right)-\lambda=0.$
Rearranging the terms to isolate the log-probability ratio:
$\log\frac{\pi(o_{i})}{\pi_{ref}(o_{i})}=\frac{N_{i}A_{i}}{\beta G\cdot\pi(o_{i})}-\left(1+\frac{\lambda}{\beta}\right).$
Then, we have:
$\log\hat{\pi}(o_{i})=\log\pi_{ref}(o_{i})+\frac{N_{i}A_{i}}{\beta G}+C,$
where $C$ is a normalization constant independent of $o_{i}$. Applying the constraint $\sum_{i}\hat{\pi}(o_{i})=1$ yields Theorem 3.2. ∎
Appendix C Distributional Update Geometry
Theorem C.1.
@@ Appendix C Distributional Update Geometry
The batch-optimal policy $\hat{\pi}$ satisfies the stationary condition $\nabla_{\pi}\hat{J}=0$, which yields $\hat{\pi}(o_{i})=\frac{1}{Z^{\prime}}\pi_{\text{ref}}(o_{i})\exp(\frac{N_{i}A_{i}}{\beta G})$ according to Theorem 3.2. Taking the log:
@@
Appendix D Proof of Theorem 3.3
@@
Appendix E Background for Section 3.3
@@
Appendix F Assumption Validation
We empirically validate Assumptions 3.1 and 3.3 on Qwen3‑0.6B. We first randomply extract 1000 pairs of questions $(q,q^{\prime})$ from GSM8K and, for each query, sample 8 responses. For $(q,o)$ and $(q^{\prime},o^{\prime})$, we compute the inner product of their logit gradients $\nabla_{\theta}f(q,o)^{\top}\nabla_{\theta}f(q^{\prime},o^{\prime})$. Across all such pairs, the mean inner product is 1467.50 with a standard deviation of 363.97, indicating that gradients for semantically related outputs are strongly and consistently positively aligned, in line with Assumption 3.1.
We then estimate the empirical kernel matrix $\mathbf{K}_{qq}=\mathbf{J}_{q}\mathbf{J}_{q}^{\top}$ for $q$. As shown in Figure 3, the diagonal entries (self‑alignment) are substantially larger than the off‑diagonal entries (cross‑alignment), corroborating the diagonal‑dominance condition in Assumption 3.3.
Figure 3: Empirical validation of diagonal dominance on Qwen3‑0.6B using GSM8K. Empirical kernel matrix $\mathbf{K}_{qq}$, where diagonal entries (self‑alignment) are larger than off‑diagonal entries (cross‑alignment).
Appendix G Proof of Theorem 3.4
Theorem G.1.
@@ Appendix F Assumption Validation
According to 3.2, we have $\mathbf{k}_{q^{\prime}}]_{s}=k((q^{\prime},o^{\prime}),(q,o^{(s)}))\approx\eta K((q,o^{\prime}),(q,o^{(s)}))$.
@@
Appendix H Proof of Corollary 3.2
Corollary H.1.
@@
Appendix I Supplementary Experiment for Section 3.4
@@
Appendix J Supplementary Material for Section 4
@@
Appendix K Supplementary Experiment for Section 5
Table 4: Comparison of AVG@8 performance in DeepSeek-R1-Distill-Qwen-1.5B trained on DAPO-Math-17K.
Method Math500 O-bench
GRPO 80.07 42.65
DAPO 82.25 43.64
IAC (Ours) 84.61 45.54
Table 5: Sensitivity of IAC to the calibration strength $\alpha$. We report AVG@8 and PASS@8 on Math500 and O-Bench for Qwen3-8B when varying $\alpha$ in IAC. When $\alpha=0$, IAC reduces to the vanilla GRPO objective. We also include a simple filtering baseline that discards the easiest and hardest 25% of training queries (by empirical success rate), retaining only the middle 50% of the difficulty spectrum.
Metric Dataset 0 0.5 1 1.5 2 Filtering-based (Bae et al., 2025)
@@
Performance comparison across different datasets and models. We further evaluate IAC on DAPO-Math-17K (Yu et al., 2025) using DeepSeek-R1-Distill-Qwen-1.5B to verify its generality. As shown in Table 4, IAC achieves the best performance, reaching 84.61% on Math500 and 45.54% on O-bench. This represents consistent improvements over GRPO (80.07% / 42.65%) and DAPO (82.25% / 43.64%), demonstrating that IAC can effectively enhance mathematical reasoning capabilities when applied to a different backbone model and training corpus.
@@
Table 5 also compares IAC against a filtering strategy that discards the easiest 25% of training queries according to their empirical success rate (Bae et al., 2025). Across both Math500 and O-Bench, filtering yields only modest gains over GRPO and is consistently weaker than IAC. This gap highlights a key limitation of hard filtering: removing easy queries entirely reduces the effective training distribution and discards useful gradient information, particularly about how to maintain correctness on already-solved patterns. IAC, in contrast, keeps all queries in the optimization loop but rescales their influence on the update. Easy queries contribute a smaller—but non-zero—signal, preventing them from overwhelming the batch partition function $Z^{\prime}(q)$, while hard queries are emphasized in a smooth, data-dependent manner. From a distributional perspective, filtering crudely prunes regions of the solution space, whereas IAC gently rebalances probability mass across difficulty levels. The latter is more effective at preserving useful behaviors on easy instances while still allocating sufficient capacity to explore and stabilize difficult, low-success regimes, which explains the consistent empirical advantage of IAC over filtering-based baselines.